# Zero-shot forecasting of chaotic systems

### v1

@@ Abstract
Time-series forecasting is a challenging task that traditionally requires specialized models custom-trained for the specific task at hand. Recently, inspired by the success of large language models, foundation models pre-trained on vast amounts of time-series data from diverse domains have emerged as a promising candidate for general-purpose time-series forecasting. The defining characteristic of these foundation models is their ability to perform zero-shot learning, that is, forecasting a new system from limited context data without explicit re-training or fine-tuning. Here, we evaluate whether the zero-shot learning paradigm extends to the challenging task of forecasting chaotic systems. Across $135$ distinct chaotic dynamical systems and $10^{8}$ timepoints, we find that foundation models produce competitive forecasts compared to custom-trained models (including NBEATS, TiDE, etc.), particularly when training data is limited. Interestingly, even after point forecasts fail, foundation models preserve the geometric and statistical properties of the chaotic attractors, demonstrating a surprisingly strong ability to capture the long-term behavior of chaotic dynamical systems. Our results highlight the promises and pitfalls of foundation models in making zero-shot forecasts of chaotic systems.
@@ 1 Introduction
Figure 1: Chaos as a benchmark for zero-shot forecasting of time series. We use the $135$ chaotic systems from the dysts dataset (Gilpin, 2021) to generate chaotic trajectories from different initial conditions. Each trajectory is used to train the baseline deep-learning models (NBEATS, TiDE, etc.) and also provided as context to the pre-trained LLM (we use Chronos, a best-in-class foundation model for time series capable of zero-shot forecasting). Both the trained baseline models and Chronos are then asked to extend the trajectory by predicting the flow of the chaotic system. We measure the quality of the predictions in terms of both short-term accuracy and long-term attractor reconstruction. Across thousands of chaotic trajectories and over $10^{8}$ data points, we found that zero-shot forecasts from Chronos can be competitive in both short-term predictions and in capturing the long-term “climate” of the dynamics.
@@
On the more theoretical side of the inquiry, an emerging direction in SciML is to understand the out-of-distribution generalization ability of different data-driven modeling frameworks (Wang et al., 2020; Kong et al., 2021, 2023; Göring et al., 2024). This parallels a long line of research that investigates the generalization ability of neural networks (Neyshabur et al., 2018; Belkin et al., 2019; Baldassi et al., 2020; Xu et al., 2020; Feng & Tu, 2021; Nakkiran et al., 2021; Power et al., 2022; Liu et al., 2022c). For example, if a model was only trained on trajectories from a limited number of initial conditions, can it effectively extrapolate the learned dynamics to a different part of the phase space and forecast from a previously unseen initial condition (that is, far from any of the training initial conditions) (Zhang & Cornelius, 2023)? Foundation models that not only generalize to new initial conditions but also to new systems could introduce novel ideas and insights into this endeavor.
@@ 2 Related work
Several recent zero-shot forecasting models are modifications of large language models, which encode time series as tokens (Xue & Salim, 2023; Ansari et al., 2024; Gruver et al., 2024; Miller et al., 2024; Liu et al., 2024). Several of these models have been shown to exhibit in-context learning at test time (Lu et al., 2024; Gao et al., 2024; Liang et al., 2024).
@@ 3 A motivating example
Figure 2: Chronos can make meaningful zero-shot forecasts of chaotic systems. Here, we use Chronos to predict the $x$ and $y$ components of the Lorenz oscillator. Since Chronos makes probabilistic forecasts, we run the model independently $20$ times and plot the mean forecast. We note that there is very little variation between different runs, implying that the uncertainty of the forecast is low. The zero-shot forecasts match remarkably well with the ground truth for both short-term prediction and long-term attractor reconstruction.
Fig. 2 shows zero-shot forecasting of the Lorenz oscillator (defined in the appendix), a well-studied chaotic dynamical system, using Chronos-200M. The only data Chronos has access to are the $512$ data points that serve as the context for the prediction (pink curve from $t=1$ to $t=512$). Because Chronos is a univariate forecast model, we separately forecast each coordinate of the attractor and reset the model state between each forecast. Forecasting chaotic systems based on partial observations (e.g., only having access to the $x$ or $y$ coordinate of the Lorenz oscillator) is an extremely difficult task in nonlinear dynamics (Ratas & Pyragas, 2024). Despite this challenge, the prediction closely tracks the ground truth for over $100$ time steps and, even after diverging from it due to chaos, remains in the vicinity of the strange attractor.
@@
However, the performance of Chronos, while impressive, can also be fragile. Keeping everything unchanged, simply starting the context trajectory from a different initial condition on the attractor can significantly degrade the accuracy of Chronos’s prediction (see Fig. 9 in the appendix). So how good is Chronos at forecasting chaotic systems, truly? More generally, is zero-shot forecasting from foundation models a promising alternative to custom-trained models when it comes to predicting chaotic systems? To answer these questions, we next perform systematic benchmarks that average over a diverse set of chaotic systems and different initial conditions.
@@ 4 Methods
4.1 Chaos as a forecasting benchmark
A natural test of modern forecasting methods comes from the field of nonlinear dynamics, which seeks to quantify the degree to which systems are predictable. A key measure is the largest Lyapunov exponent $\lambda$, an invariant property associated with every set of differential equations that quantifies the rate at which small errors accumulate. Systems that approach a periodic orbit or an equilibrium exhibit zero or negative Lyapunov exponents because different initial conditions converge to the same state. In contrast, chaotic systems exhibit positive Lyapunov exponents, implying that small changes in the initial conditions or the model parameters lead to trajectories that (at least initially) diverge exponentially. When modeling such systems, a small error will compound over a characteristic timescale, the Lyapunov time, $\tau\equiv\lambda^{-1}$, making highly-chaotic systems (those with small $\tau$) difficult to forecast.
@@ 4.1 Chaos as a forecasting benchmark
4.2 Baseline experiments
Our baseline experiment design matches prior works (Gilpin, 2021, 2023; Godahewa et al., ; Schötz et al., 2024). For each of the $135$ chaotic dynamical systems, $20$ trajectories of length $812$ are generated, each originating from a random initial condition on the attractor. This produces a set of $2700$ ($135\times 20$) multivariate time series, which have dimensionalities between $3$ and $6$ depending on the particular dynamical system. All time series are then split into training sets consisting of the first $512$ points of each time series, with the last $300$ timepoints set aside to determine final test scores. For experiments with varying context lengths, trajectories are extended backwards in time, so that the $300$ test points remain the same.
@@ 4.2 Baseline experiments
4.3 Metrics
We use four metrics to evaluate forecasting quality: Symmetric Mean Absolute Percentage Error (sMAPE).
@@ 4.3 Metrics
where $\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}$ correspond to the true test values of a time series up to a maximum forecast horizon $T$, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},...,\hat{\mathbf{x}}_{T}$ are the predictions of a forecast model at those same timepoints. Prior studies have evaluated the suitability of various error metrics in evaluating forecast accuracy (Hyndman & Koehler, 2006; Makridakis et al., 2022), including specifically on dynamical systems prediction (Gilpin, 2021, 2023), and found that sMAPE strongly correlates with other metrics (e.g. RMSE, NRMSE, MASE, Spearman correlation) while exhibiting favorable properties like a bounded range.
@@
$\text{VPT}\equiv\text{argmax}_{t_{f}}\{t_{f}|\text{sMAPE}(\mathbf{x}_{t},\hat{\mathbf{x}}_{t})<\epsilon,\;\forall t<t_{f}\}.$
@@ 5 Results
5.1 Comparing zero-shot models and fully-trained models.
@@
Historically a prediction time of 1 Lyapunov timescale has been considered prohibitive even for fully-trained forecast models. This is because both observational and modeling error compound over this timescale (Palmer, 2000; Medio & Lines, 2001). Chronos’s ability to consistently forecast up to 1 Lyapunov time, without prior training on dynamical systems, suggests the advantages of its large-scale training on diverse time series. This scale allows it to extract generic predictive features from time series, which also prove to effectively represent nonlinear dynamics. A similar concept occurs in computer vision, in which convolutional neural networks tend to learn generic Gabor-like feature extractors in early convolutional layers (Zeiler & Fergus, 2014). The ability of Chronos to generate meaningful forecasts suggests that these learned nonlinear features, coupled with high dimensionality both in the input feature space (context length) and internal model dynamics, mitigate the intrinsic chaoticity of the underlying systems. In dynamical systems theory, recent works on Koopman operators emphasize that appropriately-selected nonlinear transformations make chaotic systems appear more linear (and thus predictable) in higher-dimensional spaces (Mezić, 2013; Brunton et al., 2022). As Chronos contains tens of millions of internal weights, it has an intrinsic advantage due to its scale, which counteracts its low inductive bias when compared to forecasting models specialized in dynamical systems, such as next-generation reservoir computers.
@@
Figure 4 plots the correlation dimension accuracy (which measures long-term attractor reconstruction quality) against the VPT (which measures short-term forecast quality) for different models. NBEATS performs the best in both metrics, likely because very high pointwise accuracy necessarily leads to high global attractor quality. Generally, this trend holds across both zero-shot models and baseline models. However, within each model class a few surprises emerge: the fully-trained small Transformer, which produced relatively weak forecasts, captures the attractor shape as accurately as the zero-shot models. This observation suggests that attention-based models, which process their entire context simultaneously, have an innate advantage in capturing the long-term structure of attractors—a finding that mirrors similar results for language models (Brown, 2020). Consistent with this interpretation, we observe weak attractor reconstruction accuracy from the LSTM and NVAR models, which both operate sequentially and downweight earlier parts of their context. To ensure that these results are not a consequence of our choice of metric, we also evaluated attractor reconstruction quality using the KL Divergence between the true and forecasted attractors, and we found the same trends (see Appendix).
5.2 Accuracy and cost of in-weights versus in-context learning for varying data sizes.
We next evaluate how variation in context length affects the performance of Chronos. We vary the context length of the base Chronos model between $5$ and its maximum value of $512$ and repeat our zero-shot forecast experiments. We also select the best-performing traditional model, NBEATS, and fully train it (including cross-validation to set the lookback window) over the same points given to Chronos as context. We find that the VPT of Chronos increases monotonically with context length, even as it reaches over $17$ Lyapunov times. This timescale extends well-beyond the $\sim\!\!1$ Lyapunov timescale over which chaotic systems typically become decorrelated ($\langle\mathbf{x}(t)\mathbf{x}(t+\tau)\rangle_{t}=0$) (Shaw, 1981). This regime also exceeds the typical range of Takens’ embedding theorem, because time series are usually lifted using delay embeddings over timescales $<\tau$.
Chronos’s increasing forecasting performance stems from its ability to perform in-context learning, in which additional context points from an attractor act analogously to prompting in language models (Brown, 2020; Li et al., 2023). This mechanism underlies our earlier observation that the model’s generalization ability improves with its size. While early points in a long context are decorrelated with the predictions, they are drawn from the same underlying distribution. Chaotic attractors are ergodic, and so they have a well-defined stationary distribution of expected states known as their natural measure (Ott, 2002). As in language prediction, the early context points provide information about the measure and the conditional probabilities of successive data points (Xie et al., 2022). A similar phenomenon is observed for reservoir computers, a type of recurrent neural network in which a trainable linear layer is applied to a random transformation of the context data (Jaeger & Haas, 2004; Pathak et al., 2018). In this setting, extended context allows the reservoir to gradually synchronize with the dynamical system being learned (Lu & Bassett, 2020).
We next consider the practical question of whether the foundation model paradigm—pretraining on domain-agnostic settings, and then specializing to a task—confers computational advantages over directly training a smaller model from scratch. We measure the walltime of training and inference on a single A100 GPU node  [^1] Walltime is an imperfect measure of computational costs, because different models are specialized for different hardware (such as GPU acceleration or parallelization). Nonetheless, walltime within a given model class provides a proxy for a model’s performance in real-world settings.. We find that the computational cost of Chronos can be favorable at long context lengths when compared to NBEATS (Fig. 5B).
@@
Figure 5: Effect of context length on performance and costs. (A) The Valid Prediction Time (VPT) for base Chronos as the context length is increased, and the VPT of NBEATS tuned and trained on the same data. For these experiments, NBEATS was trained in a multivariate setting (channel-mixing), giving it a slight advantage over Chronos. (B) The single-node walltime for zero-shot forecasts with Chronos-base, compared to the training (including hyperparameter tuning) and inference costs of NBEATS. All error ranges denote standard deviations over $135$ dynamical systems, each with $20$ different initial conditions. .
@@
The inference time of Chronos is bounded by the quadratic scaling of attention layers with the context length. This limitation motivated newer architectures like Mamba (for language) and TiDE (for time series), which exhibit linear scaling. However, despite the relatively slow inference at small context windows, we find that Chronos can be very efficient when working with long context, making it a viable choice for many practical applications. In terms of the prediction horizon, Chronos exhibits the same linear scaling of cost as auto-regressive models (RC, LSTM, NVAR, etc.).

### v2

@@ Abstract
Time-series forecasting is a challenging problem that traditionally requires specialized models custom-trained for the specific task at hand. Recently, inspired by the success of large language models, foundation models pre-trained on vast amounts of time-series data from diverse domains have emerged as a promising candidate for general-purpose time-series forecasting. The defining characteristic of these foundation models is their ability to perform zero-shot learning, that is, forecasting a new system from limited context data without explicit re-training or fine-tuning. Here, we evaluate whether the zero-shot learning paradigm extends to the challenging task of forecasting chaotic systems. Across $135$ distinct chaotic dynamical systems and $10^{8}$ timepoints, we find that foundation models produce competitive forecasts compared to custom-trained models (including NBEATS, TiDE, etc.), particularly when training data is limited. Interestingly, even after point forecasts fail, large foundation models are able to preserve the geometric and statistical properties of the chaotic attractors. We attribute this success to foundation models’ ability to perform in-context learning and identify context parroting as a simple mechanism used by these models to capture the long-term behavior of chaotic dynamical systems. Our results highlight the potential of foundation models as a tool for probing nonlinear and complex systems.
@@ 1 Introduction
Figure 1: Chaos as a benchmark for zero-shot forecasting of time series. We use $135$ distinct chaotic systems to generate chaotic trajectories from $20$ different initial conditions each. Each trajectory is used to train the baseline deep-learning models (NBEATS, TiDE, etc.) and also provided as context to the pre-trained LLM (we use Chronos, a best-in-class foundation model for time series). Both the trained baseline models and Chronos are then asked to predict the trajectory into the future. We measure the quality of the predictions in terms of both short-term accuracy and long-term attractor reconstruction. Across $10^{4}$ distinct trajectories and $10^{8}$ data points, we find that zero-shot forecasts can be competitive in both short-term predictions and in capturing the long-term “climate” of the dynamics.
@@
From a theoretical standpoint, an emerging direction in SciML is to understand the out-of-distribution generalization ability of different data-driven modeling frameworks (Wang et al., 2020; Kong et al., 2021, 2023; Göring et al., 2024). This parallels a long line of research that investigates the generalization ability of neural networks (Neyshabur et al., 2018; Belkin et al., 2019; Baldassi et al., 2020; Xu et al., 2020; Feng & Tu, 2021; Nakkiran et al., 2021; Power et al., 2022; Liu et al., 2022c). For example, if a model was only trained on trajectories from a limited number of initial conditions, can it effectively extrapolate the learned dynamics to a different part of the phase space and forecast from a previously unseen initial condition (that is, far from any of the training initial conditions) (Zhang & Cornelius, 2023)? Foundation models that not only generalize to new initial conditions but also to new systems could introduce novel ideas and insights into this endeavor.
@@ 2 Related work
Several recent zero-shot forecasting models are modifications of large language models, which encode time series as tokens (Xue & Salim, 2023; Ansari et al., 2024; Gruver et al., 2024; Miller et al., 2024; Liu et al., 2024; Ekambaram et al., 2024). Several of these models have been shown to exhibit in-context learning at test time (Lu et al., 2024; Gao et al., 2024; Liang et al., 2024).
@@ 3 A motivating example
Figure 2: Zero-shot forecasts of chaotic systems. We use Chronos to predict the $x(t)$ and $y(t)$ components of the Lorenz oscillator. The zero-shot forecasts match remarkably well with the ground truth for both short-term prediction and long-term attractor reconstruction.
Figure 2 shows zero-shot forecasting of the Lorenz oscillator (defined in the appendix), a well-studied chaotic dynamical system, using Chronos-200M. The only data Chronos has access to are the $512$ data points that serve as the context for the prediction (the gray-only part of the time series). Because Chronos is a univariate forecast model, we separately forecast each coordinate of the attractor and reset the model state between each forecast. Forecasting chaotic systems based on partial observations (e.g., only having access to the $x$ or $y$ coordinate of the Lorenz oscillator) is an extremely difficult task in nonlinear dynamics (Ratas & Pyragas, 2024). Despite this challenge, the prediction closely tracks the ground truth for over $3$ Lyapunov times and, even after diverging from it due to chaos, remains in the vicinity of the strange attractor.
@@
However, the performance of Chronos, while impressive, can also be fragile. Keeping everything unchanged, simply starting the context trajectory from a different initial condition on the attractor can significantly degrade the accuracy of Chronos’s prediction (see Figs. 5, 8 and 9). So how good is Chronos at forecasting chaotic systems, truly? More generally, is zero-shot forecasting from foundation models a promising alternative to custom-trained models when it comes to predicting chaotic systems? To answer these questions, we next perform systematic benchmarks that average over a diverse set of chaotic systems and different initial conditions.
@@ 4 Methods
A chaotic systems forecasting benchmark. Nonlinear dynamics aims to quantify the degree to which physical systems are predictable. A key concept is the largest Lyapunov exponent $\lambda$, an invariant property associated with every set of differential equations that quantifies the rate at which small errors accumulate. Systems that approach a periodic orbit or an equilibrium exhibit zero or negative Lyapunov exponents because different initial conditions converge to the same state. In contrast, chaotic systems exhibit positive Lyapunov exponents, implying that small changes in the initial conditions or the model parameters lead to trajectories that (at least initially) diverge exponentially. When modeling such systems, a small error will compound over a characteristic timescale, the Lyapunov time, $\tau\equiv\lambda^{-1}$, making highly-chaotic systems (those with small $\tau$) difficult to forecast.
@@
Baseline experiments. Our baseline experiment design matches prior works (Gilpin, 2021, 2023; Godahewa et al., ; Schötz et al., 2024). For each of the $135$ chaotic dynamical systems, $20$ trajectories of length $812$ are generated, each originating from a random initial condition on the attractor. This produces a set of $2700$ ($135\times 20$) multivariate time series, which have dimensionalities between $3$ and $6$ depending on the particular dynamical system. All time series are then split into training sets consisting of the first $512$ points of each time series, with the last $300$ timepoints set aside to determine final test scores. For experiments with varying context lengths, trajectories are extended backwards in time, so that the $300$ test points remain the same.
@@
Metrics. Following prior studies (Hyndman & Koehler, 2006; Makridakis et al., 2022; Gilpin, 2021, 2023), we use four metrics to evaluate forecast quality, including Symmetric Mean Absolute Percentage Error (sMAPE).
@@
where $\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}$ correspond to the true test values of a time series up to a maximum forecast horizon $T$, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},...,\hat{\mathbf{x}}_{T}$ are the predictions of a forecast model at those same timepoints.
@@
$\text{VPT}\equiv\text{argmax}_{t_{f}}\{t_{f}|\text{sMAPE}(\mathbf{x}_{t},\hat{\mathbf{x}}_{t})<\epsilon,\;\forall t<t_{f}\}.$ (1)
@@ 5 Results
5.1 Zero-shot models are competitive with fully-trained models in short-term accuracy.
@@
Historically a prediction time of 1 Lyapunov timescale has been considered prohibitive even for fully-trained forecast models. This is because both observational and modeling error compound over this timescale (Palmer, 2000; Medio & Lines, 2001). Chronos’s ability to consistently forecast up to 1 Lyapunov time, without prior training on dynamical systems, suggests the advantages of its large-scale training on diverse time series. This scale allows it to extract generic predictive features from time series, which also prove to effectively represent nonlinear dynamics. A similar concept occurs in computer vision, in which convolutional neural networks tend to learn generic Gabor-like feature extractors in early convolutional layers (Zeiler & Fergus, 2014). The ability of Chronos to generate meaningful forecasts suggests that these learned nonlinear features, coupled with high dimensionality both in the input feature space (context length) and internal model dynamics, mitigate the intrinsic chaoticity of the underlying systems. In dynamical systems theory, recent works on Koopman operators show that appropriately-selected nonlinear transformations make chaotic systems appear more linear (and thus predictable) in higher-dimensional spaces (Mezić, 2013; Brunton et al., 2022). As Chronos contains tens of millions of internal weights, it has an intrinsic advantage due to its scale, which counteracts its low inductive bias when compared to forecasting models specialized for dynamical systems, such as next-generation reservoir computers.
5.2 Large zero-shot models excel at long-term attractor reconstruction.
@@
Figure 4 shows the correlation dimension accuracy (long-term attractor reconstruction quality) against the VPT (short-term forecast quality) for different models. NBEATS performs the best in both metrics, likely because very high pointwise accuracy necessarily leads to high global attractor quality. Generally, this trend holds across both zero-shot models and baseline models. However, within each model class a few surprises emerge: the fully-trained small Transformer, which produced relatively weak forecasts, captures the attractor shape as accurately as the zero-shot models. This observation suggests that attention-based models, which process their entire context simultaneously, have an innate advantage in capturing the long-term structure of attractors—mirroring similar results for language models (Brown, 2020). Consistent with this interpretation, we observe weak attractor reconstruction accuracy from the LSTM and NVAR models, which both operate sequentially and downweight earlier parts of their context. To ensure that these results are not a consequence of our choice of metric, we also evaluated attractor reconstruction quality using the KL Divergence between the true and forecasted attractors, and we found the same trends (see Appendix).
5.3 Zero-shot forecasts parrot motifs from their context.
We next identify a simple mechanism for zero-shot forecasting. Because Chronos is a generative time series model that learns conditional dependencies among points in its context, we directly quantify the similarity between the timepoints immediately preceding a forecast and previous intervals seen in the context. We use the highest-correlating subsequence of duration greater than $30$ timepoints ($1$ Lyapunov time in our units) as a measure of context overlap. We find that the zero-shot model’s forecasts strongly correlate with this metric over all dynamical systems, and that this dependence is more pronounced than in the best-performing fully-trained model (Fig. 5). This suggests that much of Chronos’s performance arises from its ability to parrot context sequences, underscoring our earlier observation that Chronos primarily models conditional dependencies among timepoints.
In Appendix E, we further probe this effect by showing that zero-shot performance continuously degrades as nonstationarity is introduced into the time series. Nonstationarity represents distribution shift for time series, and it disrupts the effectiveness of context-parroting as a forecast strategy because the dynamical attractor continuously and irreversibly changes. In Appendix C, we also identify a weak correlation between forecast accuracy and the first forecast point’s natural measure density (the local density of points on a dynamical system’s attractor), underscoring how rarer dynamics challenge zero-shot predictions.
Figure 5: Context parroting as a mechanism for zero-shot forecasting. (A) Better zero-shot forecasts often have initial stages that overlap with the context. The context overlap quantifies the similarity between the last $30$ points of the context and the prior points. (B) Comparison of context overlap of the zero-shot forecasts (Chronos-base) with the best performing fully-trained model (NBEATS). The zero-shot model correlates with context significantly more than the trained models across the chaotic systems dataset (matched t-test, $N=135$, $p<10^{-3}$).
5.4 Chronos performs effective in-context learning even with shuffled context.
Chronos’s forecasting performance stems from its ability to perform in-context learning, in which early context points on an attractor act analogously to prompts in language models (Brown, 2020; Li et al., 2023). This mechanism underlies our earlier observation that the model’s generalization ability improves with its size. While early points in a long context are decorrelated with the predictions, they are drawn from the same underlying distribution, and we thus hypothesize that longer contexts provide information about the distribution of attractor states, as occurs in language models (Xie et al., 2022). We test this hypothesis by randomly shuffling all length-$k$ sequences of successive timepoints in the model’s context, and then repeating our zero-shot experiments as $k$ increases (Fig. 6A). For example, if the context is $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\mathbf{x}_{4}$, then a $1$-gram shuffle would be $\mathbf{x}_{1},\mathbf{x}_{4},\mathbf{x}_{2},\mathbf{x}_{3}$ while a $2$-gram shuffle would be $\mathbf{x}_{3},\mathbf{x}_{4},\mathbf{x}_{1},\mathbf{x}_{2}$. We keep the last $k$ context timepoints the same as the original training dataset, but we ensure that the penultimate $k$ sequence differ from the unshuffled context. As a baseline, we also directly perform zero-shot forecasts using only the last $k$ context timepoints.
We find that the model’s forecast accuracy increases with the context length, but that, for sufficiently long contexts, random shuffles provide better forecasts than shorter context baselines. Earlier context points thus provide statistical information about the distribution of single timepoint values, as well as conditional probabilities of certain pairs, triplets, et cetera (Xie et al., 2022). The ergodicity of chaotic attractors implies that they have a well-defined stationary distribution of expected states $p(\mathbf{x_{t}})$, known as the natural measure (Ott, 2002). Long contexts (even when shuffled), beyond the timescale over which the states of a system become decorrelated, facilitate in-context learning of this measure. Consistent with this observation, in Appendix E, we show that non-stationary time series (in which this measure irreversibly deforms) generally lead to worse zero-shot forecasts. This process resembles the warm-up time in reservoir computers, a type of recurrent neural network used for dynamical systems forecasting (Jaeger & Haas, 2004; Pathak et al., 2018). In this setting, extended context allows the reservoir to gradually synchronize with the dynamical system being learned (Lu & Bassett, 2020).
5.5 Zero-shot forecasting is computationally efficient compared to fully-training models.
@@
Figure 6: Scaling laws with context length. (A) The forecast accuracy (VPT) of Chronos-base when given a context of length $k$ versus when given a full context (length $512$) but with all $k$-grams shuffled. (B) Comparing Chronos-base zero-shot forecasts with NBEATS fully trained on the same context. For these experiments, NBEATS was trained in a multivariate setting (channel-mixing), giving it a slight advantage over Chronos. (C) The single-node walltime for zero-shot forecasts (Chronos-base), compared to the training and inference costs of NBEATS (including hyperparameter tuning). All curves show medians and standard errors over $20$ different initial conditions from each of $135$ dynamical systems.
We next evaluate how variation in context length affects the performance of Chronos. We vary the context length of the base Chronos model between $5$ and its maximum value of $512$ and repeat our zero-shot forecast experiments. We also select the best-performing traditional model, NBEATS, and fully train it (including cross-validation to set the lookback window) over the same points given to Chronos as context. We find that the VPT of Chronos increases monotonically with context length, even as the context reaches over $17$ Lyapunov times (Fig. 6B). This timescale extends well-beyond the $\sim\!\!1$ Lyapunov timescale over which chaotic systems typically become decorrelated ($\langle\mathbf{x}(t)\mathbf{x}(t+\tau)\rangle_{t}=0$) (Shaw, 1981). This regime also exceeds the typical range of Takens’ embedding theorem, because time series are usually lifted using delay embeddings over timescales $<\tau$. Chronos’s performance therefore arises from more than just featurization and extrapolation from recent points in the context.
We next consider the practical question of whether the foundation model paradigm—pretraining in domain-agnostic settings, and then specializing to a task—confers computational advantages over directly training a smaller model from scratch. We measure the walltime of training and inference on a single A100 GPU node.  [^1] Walltime imperfectly measures computational costs, as different models are specialized for different hardware (e.g. paralleization or GPU acceleration). Nonetheless, walltime within a given model class provides a proxy for a model’s practical performance. We find that the computational cost of Chronos can be favorable at long context lengths when compared to NBEATS (Fig. 6C).
@@
The inference time of Chronos is bounded by the quadratic scaling of attention layers with the context length. This limitation motivates newer architectures like Mamba (for language) and TiDE (for time series), which exhibit linear scaling. However, despite the relatively slow inference at small context windows, we find that Chronos can be very efficient when working with long context, making it a viable choice for many practical applications. In terms of the prediction horizon, Chronos exhibits the same linear scaling of cost as auto-regressive models (RC, LSTM, NVAR, etc.).